OpenStudy (anonymous):

Tell whether each equation is linear. If so, write equation in standard form and give the values of A,B,C 2x - 8y = 16

OpenStudy (anonymous):

This equation is linear because it has no exponents. The first thing you do when trying to write a linear equation is get it into a y = mx + b form. So you want to get y alone by itself on one side. For this equation you can minus a 2x from each side. Since the left side has a 2x, the 2x - 2x will equal 0 and it is gone. Then, on the 16 side, you minus a 2x as well resulting in 16 - 2x. Now we have -8y = 16 - 2x. We don't want any numbers behind the y so what we do is divide each side by -8. Remember that if you divide any number by the same number you get 1 and multiplying 1 by any number (except 0) will give you the same number. So on the -8y side, we divide that by -8 and get 1y which equals y. On the other side we must remember that you have to divide each term (everything between negative or positive signs) by the same number we divided on the -8y side. So we divide 16 by -8 and get -2 and -2x by -8 in which we get -2x/-8 which equals (-2/-8)x, or also known as (1/4)x. We finally end up with a linear equation which is y = -2 + (1/4)x.